Article #582: Network outage affecting Peregrine1 cluster
On April 24, 2013, network engineers will be relocating fiber optics that connect the Peregrine1 cluster to infrastructure in West Lafayette. This out...
On April 24, 2013, network engineers will be relocating fiber optics that connect the Peregrine1 cluster to infrastructure in West Lafayette. This out...
Update: 8:12pm Scheduling on Carter has been resumed, and Carter is back in full production. Original Message: Beginning the morning of April 16, a nu...
Update: ITaP engineers have corrected the issue affecting the LustreC filesystem. The system is back in production. Job scheduling on Carter, Hansen a...
As of 9:00am, are seeing a problem with the LustreC scratch filesystem that serves Carter, Hansen, and Peregrine1. To prevent any more jobs from runn...
Update: As of about 11:00 am, the problem with the chilled water has been corrected, and scheduling has resumed on all RCAC clusters. Thank you for yo...
Campus chilled water serving the MATH data center is experiencing above-normal temperatures, and as a precaution, scheduling on the Coates, Rossmann,...
Update: Noon, 1/8/13 The power issue in MATH has been resolved. Power has been restored to the nodes in the Coates-A subcluster affected by the outage...
During scheduled network maintenance on network equipment connecting storage to ITaP clusters, all scheduling will be paused from 4-6pm. Running jobs...
Update: 10:00pm Tuesday As of 8:30pm Tuesday 21 August 2012, the LustreB filesystem has been returned to full service. Our storage engineers with assi...
Update - April 11, 2012 240pm At around 240pm, ITaP engineers have restored communications between the HPSS system and the tape library. Access to For...
Update : 1:45pm As of As of 1:45pm this afternoon, systems staff have completed patching the samba servers used to access storage systems. You should...
Update - 6:45 pm Tuesday, 10 April 2012 ITaP engineers have found and repaired the network issue that was affecting Coates nodes type B, C and E. Job...
Update - 9:30pm, 4/1/2012: As of about 9:30pm, Sunday, 1 April, ITaP systems staff have returned Hansen to production status, and job scheduling is re...
Due to a network issue, the server running the PBS software for Rossmann is unavailable. While the server is unavailable, attempts to use PBS commands...
At approximately 10:50pm, Thursday, March 15, the power distribution to large portions of the Rossmann cluster failed. These feeds also power the logi...
Update: As of 9:45pm, Lustre is back in production and scheduling has resumed on Hansen. Original Notice: As of approximately 8:00pm February 7, an is...
This morning, the PBS system on Coates developed an issue with the storage holding its internal state.While systems engineers are working on recoverin...
Update - 1/9/2012 The repairs to the ADIC tape library have been completed and Fortress' tape functionality is back in operation. Update - 1/6/2012 Fo...
Update The error condition on the Lustre filesystem has been cleared, and Hansen is back in production and accepting new jobs. Jobs already running sh...
Update 12/2/11 (4:15pm) The tape robot has been returned to service and Fortress is back in production. Please contact us at rcac-help@purdue.edu if y...